OpenStudy (anonymous):

A number of people were asked how many days they would wait to seek medical treatment if they were suffering pain that interfered with their ability to work. The results are presented in the following table. Consider this collection of people to be a population. Let X be the number of days for a person sampled at random from this population. Find the probability that a person would wait more than 2 days.

OpenStudy (anonymous):

see screen shot for table:

OpenStudy (kropot72):

Adding the values in the frequency column, we get 862. Therefore there were 862 people in the survey. Waiting time (days) Probability 3 191/862 4 19/862 5 10/862 The probability that a randomly selected person from the population would wait more than 2 days is given by: \[\large P(more\ than\ 2\ days)=\frac{191}{862}+\frac{19}{862}+\frac{10}{862}=you\ can\ calculate\]